OpenStudy (anonymous):

for 2x + 5y = -10 find the x-intercept find the y-intercept find one other point if x = 5 graph the line

OpenStudy (anonymous):

Okay one second

OpenStudy (anonymous):

both the x and y intercepts are 0

OpenStudy (anonymous):

How did you get that

OpenStudy (anonymous):

you have to isolate the y first

OpenStudy (anonymous):

Do you know how to do that?

OpenStudy (anonymous):

2x + 5y = -10 /5 /5 =2x+y=-2

OpenStudy (anonymous):

2x+y=-2 -2x -2x y=-2x-2

OpenStudy (anonymous):

WOAH oh my god i was so wrong haha sorry my calculator was on the wrong setting

OpenStudy (anonymous):

x inercpet is -5 y intercept is -2

OpenStudy (anonymous):

Those are the correct intercepts

OpenStudy (anonymous):

what about finding another point if x = 5

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

How would you graph that?

OpenStudy (anonymous):

Do you have a graphing calculator?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

well ill give you a few points to graph and you just connect the dots

OpenStudy (anonymous):

5,-4 0,-2 -5,0

OpenStudy (anonymous):

Got it?